# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v2

@@ 1 Introduction
Infectious disease outbreaks consistently challenge public health systems, affecting both individual well-being and economic stability Nicola et al. (2020). Effective management of these outbreaks hinges on accurate epidemic forecasting, which involves predicting future incidences like infection cases and hospitalizations Liu et al. (2024b); Wan et al. (2024); Adhikari et al. (2019). Over the years, various models have been developed to address this need. These include mechanistic models like SIR Cooper et al. (2020) and statistical models like ARIMA Sahai et al. (2020); Kontopoulou et al. (2023), as well as advanced machine learning methods such as LSTM and GRU Shahid et al. (2020), which have proven instrumental in forecasting disease spread and supporting informed public health decision-making.
@@
Despite the advancements, current models are typically trained for specific diseases within particular geographic regions, limiting their ability to integrate insights from diverse sources spanning multiple pathogens and spatiotemporal contexts. This narrow focus can impede a comprehensive understanding of disease dynamics and the design of effective outbreak responses, especially during novel or emergent outbreaks when observations are typically scarce. Given the extensive and diverse outbreak data collected over decades and across various geographies, pre-training on such a broad dataset could potentially enable the development of more generalizable models with greater applicability and adaptability across different pathogens and contexts. This raises an important question: Can we leverage lessons from diverse historical disease time series to develop a generalized model that enhances epidemic forecasting accuracy?
@@
To address the above question, we draw inspiration from the success of large pre-trained transformer-based models Zhao et al. (2023) and develop a pre-trained epidemic forecasting model using extensive disease time series data to distill generalizable knowledge across pathogens and contexts. The pre-trained model can be subsequently fine-tuned for specific diseases or geographical regions. While it is possible to adapt general time series foundation models Liang et al. (2024); Ma et al. (2024) to epidemic forecasting, their pre-trained corpus mostly consists of non-epidemic data, which may not accurately capture epidemic dynamics and infection trajectories, potentially degrading forecasting accuracy. Although an early effort has been made in epidemic pre-training Kamarthi & Prakash (2023), it overlooks critical external factors such as temperature, elevation, and public health policies and interventions – factors are known to influence the dynamics of disease spread in space and time Lau et al. (2020b) – potentially yielding suboptimal performance. For instance, dengue infection spread may exhibit distinct dynamics in different geographical regions due to variations in temperature and humidity Chen & Hsieh (2012). Without accounting for these external factors, models risk failing to capture their complex interplay with pathogens and producing inaccurate forecasts. Throughout this paper, we refer to these external factors as environments.
@@
Nevertheless, the need to robustly and effectively account for the environment further intensifies the challenge of developing an epidemic pre-training framework that is generalizable across varying pathogens and contexts. A major obstacle is the shift in the temporal distribution of infection trajectories between training and test datasets, often driven by the changes in the environment. Insufficient consideration of such distribution shifts can obscure the relationship between historical infection data and future predictions (for a detailed discussion, see Appendix A.8), compromising a model’s ability to make accurate forecasts. As such, it is crucial to disentangle the influence of changing environments from other more intrinsic factors (e.g., a pathogen’s infection rate) affecting disease transmission dynamics. Yet, exact and explicit mechanisms by which the environment influences the disease dynamics of a particular pathogen are often not fully understood, which necessitates a sophisticated modeling approach to identify and separate these latent environmental influences.
@@
Figure 2: Structural causal model for epidemic foresting, where $Z$ refers to the environment states, and $X_{s}$ and $X_{c}$ refer to the spurious and causal factors of the input.
Our Solution. To integrate insights from extensive historical diseases and effectively model environmental factors, we propose Covariate-Adjusted Pretraining for Epidemic forecasting (CAPE) to capture the universal patterns of disease dynamics, as shown in Figure 2. Our approach addresses the challenges of optimizing the model with limited observations of a single disease infection trajectory and the complex influence of the environment by combining a pre-training framework with explicit environment modeling. Drawing on principles from causal analysis and covariate adjustment Runge et al. (2023), CAPE aims to estimate the latent environments and control for their influences for epidemic forecasting. Specifically, during the pre-training phase, CAPE utilizes environment-aware self-supervised learning, including random masking (Figure 2(c)) and hierarchical environment contrasting (Figure 2(b)), to enhance its understanding of the disease dynamics and environmental influence. Furthermore, an environment estimator is introduced, which estimates dynamic environments based on latent environment representations learned during pre-training using Expectation-Maximization algorithm. Our contributions can be summarized as follows:
  1. We propose a novel epidemic pre-training framework, namely CAPE, that learns representations of environments and performs covariate adjustment on the input epidemic time series data, which aims to disentangle the inherited disease dynamics from the environment.
@@
  2. We assemble a diverse collection of epidemic time series datasets from various diseases and regions, serving as a crucial testbed for evaluating pre-trained epidemic forecasting models. This allows for extensive testing across multiple scenarios, including few-shot, zero-shot, cross-location, and cross-disease evaluations.
@@
  3. We demonstrate the effectiveness of pre-training on epidemic datasets, showcasing superior performance across various downstream datasets and settings. Notably, CAPE surpasses the best baseline by an average of 9.9% in the full-shot setting and 18.1% in the zero-shot setting across all tested downstream datasets.
@@
  4. We provide an in-depth analysis of how pre-training and environment estimation affect downstream performance and mitigate the impact of distribution shifts.
@@ 2 Related Work and Problem Definition
Epidemic Forecasting Models. Traditionally, epidemic forecasting employs models like ARIMA Sahai et al. (2020), SEIR He et al. (2020), and VAR Shang et al. (2021). ARIMA predicts infections by analyzing past data and errors, SEIR models population transitions using differential equations, and VAR captures linear inter-dependencies by modeling each variable based on past values. Recently, deep learning models—categorized into RNN-based, MLP-based, and transformer-based—have surpassed these methods. RNN-based models like LSTM Wang et al. (2020) and GRU Natarajan et al. (2023) use gating mechanisms to manage information flow. MLP-based models use linear layers Zeng et al. (2023) or multi-layer perceptrons Borghi et al. (2021); Madden et al. (2024) for efficient data-to-prediction mapping. Transformer-based models Wu et al. (2021); Zhou et al. (2021, 2022) apply self-attention to encode time series and generate predictions via a decoder. However, these models are limited in that they typically utilize data from only one type of disease without considering valuable insights and patterns from diverse disease datasets.
@@
Pre-trained Time Series Models. To enhance performance and enable few-shot or zero-shot capabilities, transformer-based models often employ pre-training on large datasets, which typically use masked data reconstruction Zerveas et al. (2021); Rasul et al. (2023) or promote alignment across different contexts Fraikin et al. (2023); Zhang et al. (2022); Yue et al. (2022). For example, PatchTST Nie et al. (2022) segments time series into patches, masks some, and reconstructs the masked segments. Larger foundational models like MOMENT Goswami et al. (2024) aim to excel in multiple tasks (e.g., forecasting, imputation, classification) but require substantial data and computational resources. In the epidemic context, Kamarthi et al. Kamarthi & Prakash (2023) pre-trained on various diseases, improving downstream performance and highlighting pre-training’s potential in epidemic forecasting. Nevertheless, all these models overlook the influence of the environment, and zero-shot ability in epidemic forecasting, along with the factors affecting the pre-training process, remain unanswered. In this study, we introduce environment modeling and conduct a thorough analysis of these questions.
@@
Problem Definition. In this study, we adopt a univariate setting: Given a historical time series input: $\mathbf{x}\in\mathbb{R}^{T\times 1}$, where $T$ is the size of lookback window, the goal of epidemic forecasting is to map $\mathbf{x}$ into target trajectories (e.g. infection rates): $\mathbf{y}\in\mathbb{R}^{h}$, where $h$ denotes the size of the forecast horizon. We define $X$ and $Y$ as the random variables of input $\mathbf{x}$ and target $\mathbf{y}$ respectively. During pre-training, a representation function $g_{\theta}:\mathbb{R}^{T\times 1}\to\mathbb{R}^{T\times d}$, where $d$ denotes the dimension of the latent space and $\theta$ being the parameter of the model, extracts universal properties from a large collection of epidemic time series datasets $\mathcal{D}_{\text{pre}}=\{D_{1}^{\prime},D_{2}^{\prime},\dots,D_{S}^{\prime}\}$. Then, a set of self-supervised tasks $\mathcal{T}_{\text{pre}}=\{\mathcal{T}_{i}\}_{i=1}^{R}$ is defined, where each task $\mathcal{T}_{i}$ transforms a sample $\mathbf{x}\sim\mathcal{D}_{\text{pre}}$ into a pair of new input and label: $(\mathbf{\tilde{x}},\mathbf{\tilde{y}})$, and optimizes a loss $\mathcal{L}_{\mathcal{T}_{i}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{pre}}}[\ell_{\mathcal{T}_{i}}(h_{\psi}(g_{\theta}(\mathbf{\tilde{x}})),\mathbf{\tilde{y}})]$, with $\ell_{\mathcal{T}_{i}}$ being the task-specific metric and $h_{\psi}$ the task-specific head.
@@ 3 Proposed Method
3.1 Model Design
3.1.1 Causal Analysis for Epidemic Forecasting
@@
As environments influence both historical infection patterns and future disease spread, we draw inspiration from causal inference Zhou et al. (2023); Jiao et al. (2024) and introduce a Structural Causal Model where we treat the environment $Z$ as a confounder that influences both the independent variable (e.g., historical data $X$) and the dependent variable (e.g., future infections $Y$). Furthermore, we adopt a causal decomposition approach Mao et al. (2022) that separates $X$ into two components (Figure 2): (1) a spurious factor $X_{s}$ that is environment-dependent, and (2) a causal factor $X_{c}$ that remains environment-independent. Both factors influence the target $Y$, with $X_{s}$ reflecting the impact of environment $Z$. Since epidemic dynamics are driven by a finite set of critical factors, such as public health policies, we model $Z$ with the following assumption:
Assumption 3.1.
@@
The environment variable $Z$ follows a categorical distribution $p(Z)$ and takes on one of $K$ discrete environmental states, denoted as $z_{k}$. Each state $z_{k}$ is associated with a unique latent representation $\mathbf{e}_{k}\in\mathbb{R}^{h_{e}}$, capturing the unique features specific to that environment.
@@
In constructing a predictive model for input $\mathbf{x}$, we define $\hat{Y}$ as the predicted time series $\hat{\mathbf{y}}$ and model the predictive distribution $p_{\Theta}(\hat{Y}|X)$ using $f_{\Theta}(\mathbf{x})=h_{\psi}(g_{\theta}(\mathbf{x}))$, where $\Theta=\{\theta,\psi\}$. Training typically involves maximizing the log-likelihood of $p_{\Theta}(\hat{Y}|X)$, which in practice translates to minimizing the errors over the pre-training dataset $\mathcal{D}_{\text{pre}}$:
$\Theta^{*}=\arg\min_{\Theta}-\frac{1}{|\mathcal{D}_{\text{pre}}|}\sum\nolimits_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{\text{pre}}}\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}.$ (1)
As the environment $Z$ impacts the distribution of the observed data through $p(X,Y|Z)=p(X|Z)p(Y|X,Z)$, we formulate the following objective:
$\small\begin{split}\Theta^{*}=&\arg\min\nolimits_{\Theta}\mathbb{E}_{p(Z)}[\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p(Y,X|Z)}[\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}]].\end{split}$ (2)
@@
The above equation suggests that the optimal $\Theta^{*}$ depends on the environment distribution $p(Z)$. If we simply maximize the likelihood $p_{\Theta}(\hat{Y}|X)$, the confounding effect of $Z$ on $X$ and $Y$ will mislead the model to capture the shortcut predictive relation between the input and the target trajectories, which necessitates explicit modeling of the environment during pre-training. Given that input infection trajectories inherently reflect the influence of the environment, it is crucial to develop mechanisms that disentangle the correlations between infection trajectories and environmental factors.
@@
However, obtaining detailed environmental information, or ${\bf e}_{k}$, can be challenging due to variability in data availability and quality. To address this, we resort to a data-driven approach that treats ${\bf e}_{k}$ as learnable parameters and thus allows us to dynamically infer the environmental distribution directly from the observed data. Specifically, we implement an environment estimator $q_{\phi}(Z|X)$ that infers the probability of environment states based on historical inputs together with the latent representations of each state. Then, we derive a variational lower bound (see Appendix A.1):
$\log{p_{\Theta}(\hat{Y}|do(X))}\geq$ (4)
$\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]-\operatorname{KL}\left(q_{\phi}(Z|X)\parallel p(Z)\right),$

### v3

@@ 1 Introduction
Infectious disease outbreaks consistently challenge public health systems, affecting both individual well-being and economic stability (Nicola et al., 2020). Effective management of these outbreaks hinges on accurate epidemic forecasting, which involves predicting future incidences like infection cases and hospitalizations (Liu et al., 2024b; Wan et al., 2024; Adhikari et al., 2019). Over the years, various models have been developed to address this need. These include mechanistic models like SIR (Cooper et al., 2020) and statistical models like ARIMA (Sahai et al., 2020; Kontopoulou et al., 2023), as well as advanced machine learning methods such as LSTM and GRU (Shahid et al., 2020), which have proven instrumental in forecasting disease spread and supporting informed public health decision-making.
@@
Despite the advancements, current models are typically trained for specific diseases within particular geographic regions, limiting their ability to integrate insights from diverse sources spanning multiple pathogens and spatial contexts. This narrow focus can impede a comprehensive understanding of disease dynamics and the design of effective outbreak responses, especially during novel or emergent outbreaks when observations are typically scarce. Taking Influenza, Measles, and Dengue as examples (see dataset details in Appendix A.4), the observed performance gap in Figure 1 between a simple ARIMA model and more complex deep learning models, especially Transformer-based models, highlights a key limitation: more complex models often struggle in epidemic forecasting when prior knowledge is unavailable and observational data is scarce. Nevertheless, given the extensive and diverse outbreak data collected over decades and across various geographies, pre-training on such a broad dataset could potentially enable the development of more generalizable models with greater applicability and adaptability across different pathogens and contexts. This raises an important question: Can we leverage lessons from diverse historical disease time series to enhance epidemic forecasting accuracy for more complex transformer-based models?
@@
To address the above question, we draw inspiration from the success of large pre-trained transformer-based models (Zhao et al., 2023) and develop a pre-trained epidemic forecasting model using extensive disease time series data to distill generalizable knowledge across pathogens and contexts. The pre-trained model can be subsequently fine-tuned for specific diseases or geographical regions. While it is possible to adapt general time series foundation models (Liang et al., 2024; Ma et al., 2024) to epidemic forecasting, their pre-trained corpus mostly consists of non-epidemic data, which may not accurately capture epidemic dynamics and infection trajectories, potentially degrading forecasting accuracy.
@@
Furthermore, although an early effort has been made in epidemic pre-training (Kamarthi & Prakash, 2023), it overlooks critical external factors, which we refer to as environment throughout this paper. These factors such as temperature, humidity, public health policies, and interventions are well-known to influence the dynamics of disease spread in space and time (Lau et al., 2020b; Chen & Hsieh, 2012). As the environment influences both historical infection trajectories and future disease spread, it can be seen as an important confounder from the perspective of causal inference (Zhou et al., 2023; Jiao et al., 2024), which hinders the learning of true causal relations between the historical infections and future outbreaks, as shown in the structured causal model in Figure 2. Driven by the changes in the environment, a major obstacle arises as the shift in the distribution of infection trajectories between training and test datasets  (Wu et al., 2024). Insufficient consideration of such distribution shifts can obscure the relationship between historical infection data and future predictions (for a detailed discussion, see Appendix A.15), compromising a model’s ability to make accurate forecasts. As such, it is crucial to disentangle the influence of changing environments from other more intrinsic factors (e.g., a pathogen’s infection rate) affecting disease transmission dynamics. Yet, exact and explicit mechanisms by which the environment influences the disease dynamics of a particular pathogen are often not fully understood or observed, which necessitates an implicit modeling approach to identify and separate these latent environmental influences.
Figure 1: Performance gap between the ARIMA model and various non-pretrained deep learning models across selected datasets and horizons. Positive values indicate that the deep learning models underperform relative to ARIMA, highlighting the limitations of complex architectures without pretraining or prior knowledge.
@@
Figure 2: Structural causal model for epidemic forecasting, where $Z$ refers to the environment as a confounder. $X_{sp}$ and $X_{ca}$ are the spurious and causal factors of the input.
Our Solution. We propose Covariate-Adjusted Pretraining for Epidemic forecasting (CAPE) to capture universal, environment-invariant disease patterns from historical data (Figure 3). CAPE tackles data scarcity and distribution shift by integrating pre-training with implicit environment modeling, grounded in causal analysis. Our contributions include:
  1. As one of the first epidemic pre-trained models, we propose a generalizable pre-training framework that adjusts for covariates in epidemic time series data, disentangling intrinsic disease dynamics from environmental influences and yielding environment-invariant representations.
@@
  2. We create a comprehensive testbed for evaluating pre-trained epidemic models by assembling epidemic time series datasets across various diseases and regions. This enables thorough assessment across various scenarios, including few-shot, zero-shot, cross-location, and cross-disease settings.
@@
  3. We demonstrate the potential of epidemic pre-training, showcasing superior performance across downstream datasets and settings. Notably, CAPE surpasses the best baseline by an average of 9.9% in the full-shot setting and 18.1% in the zero-shot setting across all tested downstream datasets.
@@
  4. We provide an in-depth analysis of how pre-training and environment modeling affect downstream performance and mitigate the impact of distribution shifts.
@@ 2 Related Work and Problem Definition
Epidemic Forecasting Models. Traditionally, epidemic forecasting employs models like ARIMA (Sahai et al., 2020), SEIR (He et al., 2020), and VAR (Shang et al., 2021). ARIMA predicts infections by analyzing past data and errors, SEIR models population transitions using differential equations, and VAR captures linear inter-dependencies by modeling each variable based on past values. Recently, deep learning models—categorized into RNN-based, MLP-based, and transformer-based—have surpassed these methods. RNN-based models like LSTM (Wang et al., 2020) and GRU (Natarajan et al., 2023) use gating mechanisms to manage information flow. MLP-based models use linear layers (Zeng et al., 2023) or multi-layer perceptrons (Borghi et al., 2021; Madden et al., 2024) for efficient data-to-prediction mapping. Transformer-based models (Wu et al., 2021; Zhou et al., 2021, 2022) apply self-attention to encode time series and generate predictions via a decoder. However, these models are limited as they typically utilize data from only one type of disease without considering valuable insights from diverse disease datasets.
@@
Pre-trained Time Series Models. To enable few-shot or zero-shot capabilities, transformer-based models often employ pre-training on large datasets, which typically use masked data reconstruction (Zerveas et al., 2021; Rasul et al., 2023) or promote alignment across different contexts (Fraikin et al., 2023; Zhang et al., 2022; Yue et al., 2022). For example, PatchTST (Nie et al., 2022) segments time series into patches, masks some, and reconstructs the masked segments. Larger foundational models like MOMENT (Goswami et al., 2024) aim to excel in multiple tasks (e.g., forecasting, imputation, classification) but require substantial data and computational resources. In epidemic contexts, Kamarthi et al. (Kamarthi & Prakash, 2023) pre-train a model on various diseases, improving downstream performance and highlighting pre-training’s potential in epidemic forecasting. However, these works overlook environmental influence and zero-shot ability in epidemic forecasting and lack a deep analysis of how pre-training materials impact downstream performance. In this study, we introduce environment modeling and conduct a thorough analysis of these questions (see A.5 for more discussions).
@@
Problem Definition. Given a historical time series input: $\mathbf{x}\in\mathbb{R}^{T\times 1}$, where $T$ is the size of lookback window, the goal of epidemic forecasting is to map $\mathbf{x}$ into target trajectories (e.g. infection rates): $\mathbf{y}\in\mathbb{R}^{h}$, where $h$ denotes the size of the forecast horizon. We define $X$ and $Y$ as the random variables of input and target, respectively. During pre-training, a representation function $g_{\theta}:\mathbb{R}^{T\times 1}\to\mathbb{R}^{T\times d}$, where $d$ denotes the dimension of the latent space and $\theta$ being the parameter of the model, extracts universal properties from a large collection of epidemic time series datasets $\mathcal{D}_{\text{pre}}=\{D_{1}^{\prime},D_{2}^{\prime},\dots,D_{S}^{\prime}\}$. Then, a set of self-supervised tasks $\mathcal{T}_{\text{pre}}=\{\mathcal{T}_{i}\}_{i=1}^{R}$ is defined, and each $\mathcal{T}_{i}$ transforms a sample $\mathbf{x}\sim\mathcal{D}_{\text{pre}}$ into a new input-label pair: $(\mathbf{\tilde{x}},\mathbf{\tilde{y}})$, and optimizes a loss $\mathcal{L}_{\mathcal{T}_{i}}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{pre}}}[\ell_{\mathcal{T}_{i}}(h_{\psi}(g_{\theta}(\mathbf{\tilde{x}})),\mathbf{\tilde{y}})]$, with $\ell_{\mathcal{T}_{i}}$ being the task-specific metric and $h_{\psi}$ the task-specific head.
@@ 3 Proposed Method
Pre-training models for epidemic forecasting using data from diverse contexts presents two primary challenges. First, the inherent heterogeneity stemming from varied environmental conditions (e.g., public health interventions, seasonal factors, population behaviors) necessitates a model architecture capable of accounting for these influences. Second, the pre-training objectives must be designed to foster the learning of robust representations that capture these intrinsic dynamics and generalize effectively across diverse epidemiological scenarios, including novel or data-scarce situations. We will elaborate on our solutions in the following subsections to address these challenges.
3.1 Causal Analysis for Epidemic Forecasting
@@
In epidemic forecasting, environmental factors, such as seasonal changes and population mobility, are known to confound the relationship between historical infection patterns and future disease spread. To address this critical challenge in understanding true transmission dynamics, we draw inspiration from causal inference (Zhou et al., 2023; Jiao et al., 2024) and introduce a Structural Causal Model (Figure 2) where we treat the environment $Z$ as an unobserved confounder that influences both the independent variable (e.g., historical infections $X$) and the dependent variable (e.g., future infections $Y$). Furthermore, based on the causal decomposition assumption (Mao et al., 2022), we know $X$ is composed of (1) a spurious factor $X_{sp}$ that is environment-dependent, and (2) a causal factor $X_{ca}$ representing the underlying, more intrinsic disease transmission dynamics (e.g., pathogen infectivity) that remains environment-independent. While $X$ as a whole is observed, only $X_{ca}$ has a causal relation with target $Y$. Here, we assume that a finite set of critical factors drives the epidemic dynamics:
@@
Assumption 3.1. The environment variable $Z$ follows a categorical distribution $p(Z)$ and takes on one of $K$ discrete environmental states, denoted as $z_{k}$. Each state $z_{k}$ is associated with a unique latent representation $\mathbf{e}_{k}\in\mathbb{R}^{h_{e}}$, capturing the unique features specific to that environment.
@@
In constructing a predictive model for input $\mathbf{x}$, we define $\hat{Y}$ as the predicted time series $\hat{\mathbf{y}}$ and model the predictive distribution $p_{\Theta}(\hat{Y}|X)$ using $f_{\Theta}(\mathbf{x})=h_{\psi}(g_{\theta}(\mathbf{x}))$, where $\Theta=\{\theta,\psi\}$. Training typically involves maximizing the log-likelihood of $p_{\Theta}(\hat{Y}|X)$, which in practice translates to minimizing the errors over the pre-training dataset $\mathcal{D}_{\text{pre}}$: $\Theta^{*}=\arg\min_{\Theta}-\frac{1}{|\mathcal{D}_{\text{pre}}|}\sum\nolimits_{(\mathbf{x},\mathbf{y})\in\mathcal{D}_{\text{pre}}}\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}$. As the environment $Z$ impacts the distribution of the observed data through $p(X,Y|Z)=p(X|Z)p(Y|X,Z)$, we formulate the following objective:
$\begin{split}\Theta^{*}=&\arg\min\nolimits_{\Theta}-\mathbb{E}_{p(Z)}[\mathbb{E}_{(\mathbf{x},\mathbf{y})\sim p(Y,X|Z)}[\|\mathbf{y}-f_{\Theta}(\mathbf{x}))\|^{2}]].\end{split}$ (1)
@@
The above equation suggests that the optimal $\Theta^{*}$ depends on the environment distribution $p(Z)$. If we simply maximize the likelihood $p_{\Theta}(\hat{Y}|X)$, the confounding effect of $Z$ on $X$ and $Y$ will mislead the model to capture the shortcut predictive relation between the input and the target trajectories, which necessitates implicit modeling of the environment during pre-training. Given that input infection trajectories inherently reflect the influence of the environment, it is crucial to develop mechanisms that disentangle the correlations between infection trajectories and environmental factors.
@@
However, obtaining detailed environmental information, or ${\bf e}_{k}$, can be challenging due to variability in data availability and quality. To address this, we resort to a data-driven approach that treats ${\bf e}_{k}$ as learnable parameters and thus allows us to dynamically infer the environmental distribution directly from the observed data. Specifically, we implement an environment estimator $q_{\phi}(Z|X)$ that infers the probability of environment states based on historical inputs together with the latent representations of each state. Then, a variational lower bound is derived (Wu et al., 2024) (see Appendix A.3):
$\log{p_{\Theta}(\hat{Y}|do(X))}\geq$ $\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]$ (3)
$-\operatorname{KL}\left(q_{\phi}(Z|X)\parallel p(Z)\right),$